OpenStudy (anonymous):

What is the solution to the equation? 3(x – 5) = 8x – 30 A. x = 3 B. x = 5 C. x = –3 D. x = –5

OpenStudy (anonymous):

Distribute 3 3 * x = ? 3 * -5 = ?

OpenStudy (anonymous):

what do you think it is first?

OpenStudy (anonymous):

distribute 3 x "x" and 3x-5

OpenStudy (anonymous):

Distributing gives us: 3x - 15 = 8x - 30 Add 15 to both sides -30 + 15 = ?

OpenStudy (anonymous):

= 15

OpenStudy (anonymous):

Yes, that gives us: 3x = 8x + 15 Now we subtract 8x to both sides, what's 3x - 8x?

OpenStudy (anonymous):

-5x?

OpenStudy (anonymous):

*silent claps*

OpenStudy (anonymous):

Can you help with one more?

OpenStudy (anonymous):

Yep, now we have: -5x = 15 Now we divide -5 to both sides, what's 15 / -5

OpenStudy (anonymous):

And yes, I can

OpenStudy (anonymous):

What is the solution to the inequality? 4x + 7 > –13 A. x < –5 B. x > –5 C. x < 5 D. x > 5

OpenStudy (anonymous):

hang on

OpenStudy (anonymous):

One question per post, close this one and open a new one

OpenStudy (anonymous):

\(\bbox [10pt, lime, border:5pt solid black]{\Huge\cal\color{red}\diamondsuit\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\diamondsuit}}}\) You can give medals by clicking 'Best Response'.